### v1

# Building Age Estimation: A New Multi-Modal Benchmark Dataset and Community Challenge

Abstract
Estimating the construction year of buildings is of great importance for sustainability. Sustainable buildings minimize energy consumption and are a key part of responsible and sustainable urban planning and development to effectively combat climate change. By using Artificial Intelligence (AI) and recently proposed Transformer models, we are able to estimate the construction epoch of buildings from a multi-modal dataset. In this paper, we introduce a new benchmark multi-modal dataset, i.e. the Map your City Dataset (MyCD), containing top-view Very High Resolution (VHR) images, Earth Observation (EO) multi-spectral data from the Copernicus Sentinel-2 satellite constellation, and street-view images in many different cities in Europe, co-localized with respect to the building under study and labelled with the construction epoch. We assess EO generalization performance on new/ previously unseen cities that have been held-out from training and appear only during inference. In this work, we present the community-based data challenge we organized based on MyCD. The AI4EO Challenge MapYourCity was opened in 2024 for 4 months. We present the Top-4 performing models, and the evaluation results. During inference, the performance of the models using both all three input modalities and only the two top-view modalities, i.e. without the street-view images, is examined. The evaluation results show that the models are effective and can achieve good performance on this difficult real-world task of estimating the age of buildings, even on previously unseen cities, as well as even using only the two top-view modalities (i.e. VHR and Sentinel-2) during inference.
I Introduction
Estimation of the age of buildings in cities is important for sustainability, urban planning, and structural safety purposes. Sustainable buildings minimize energy consumption and are a key part of responsible and sustainable urban planning and development that seeks to effectively combat climate change, as well as rising temperatures. With increase in urbanization and Earth Sciences research into urban areas, the need to estimate the energy efficiency of buildings, so as to meet important sustainability goals, has never been greater. The age of buildings plays an important role in energy modelling, as well as in urban policies. The age of the building can be used as a proxy for the energy efficiency. However, data for the construction period of buildings is not currently available globally in a consistent and homogeneous manner. To this end, as a start, in this study, we focus on European cities.
In this real-world application study for sustainability and climate, the main aim is to develop models that automatically estimate the construction year of buildings with improved performance using Artificial Intelligence (AI) techniques and cross-view datasets based on street-view images and Earth Observation (EO) datasets. The goals of this paper are to present the new multi-modal dataset, examine example images from this street-view and EO dataset, present the community data challenge that we organized, and present the evaluation results of the top-performing models from the challenge. We introduce a new benchmark dataset for a real-world multi-modal application study. It consists of three modalities, including street-view ground images, top-view satellite Very High Resolution (VHR) images at 50 cm resolution, and Copernicus Sentinel-2 multi-spectral satellite data. All the modalities are co-localized with respect to the specific building under study, as well as its label of the construction epoch. The problem is formulated as a classification task with seven classes for the construction epoch, ranging from 1900 to nowadays.
This paper evaluates the EO generalization performance of the models on cities that are not included in the training dataset, i.e. on new/ previously unseen cities. We examine the performance of the models when both training and testing are performed with all 3 modalities. There is benefit in using the top-view VHR images (and to a lesser extent, Sentinel-2 data), in addition to the street-view images. In this work, we also examine the performance of the models when training is performed with all the three modalities, and testing is performed without the street-view images, i.e. when using only two out of the three encoders (i.e. top-view VHR and Sentinel-2 images). These results are compared to those when training and testing are performed using all the three modalities.
Figure 1: Left: Example street-view image from the multi-modal dataset. Right: Example top-view satellite VHR image.
Main contributions. Our contributions are that this work introduces a novel multi-modal dataset, presents the challenge we organized to estimate the age of buildings in different cities in Europe using this dataset, examines the performance evaluation results of the models from the challenge, and shows the potential of effectively combining and fusing street-view, top-view satellite and remotely sensed Sentinel-2 EO data for accurately predicting the construction period of buildings. In this paper, we present and compare the performance of the models from the challenge that have been trained and tested on the new multi-modal street-view and EO dataset.
II Data and Task
II-A The multi-modal street-view and EO dataset MyCD
We present the multi-modal dataset and examine example images from this street-view and EO dataset. The dataset MyCD encompasses the three modalities of: a) Street-view images, b) Top-view satellite VHR RGB images at 50 cm resolution, and c) Multi-spectral 12-band S-2 data. These input modalities have very different characteristics (number of spectral bands, spatial resolution, and building perspectives from top and street-view). In the dataset, the three modalities are co-localized with respect to the specific building under study. In Fig. 1, we show two example images for the modalities of street-view ground and top-view satellite VHR, respectively.
Figure 2: Example top-view satellite VHR image and the corresponding geo-localized S-2 image (here, RGB bands for visualization purposes, where the dataset includes all the $12$ spectral bands, i.e. L2A) from the multi-modal dataset MyCD.
In Fig. 2, an example top-view VHR image and the corresponding S-2 image, both from the new multi-modal dataset MyCD, are shown, together with the building location in the centre, as a blue dot in the image. We observe the geospatial overlay of the two top-view image modalities and that the two top-view modalities are centered at the building under study.
In Fig. 3, additional examples of street-view and top-view images from the new multi-modal dataset MyCD are shown.
The dataset MyCD comprises cross-view image data from $6$ countries in Europe, and more specifically from $19$ cities, for $7$ different building construction epoch classes extracted from the EUBUCCO database. For the training set data, $15$ cities have been chosen, while for the test set data, the remaining $4$ cities have been used. The building construction epoch classes are $7$, i.e. labelled from 00 to $6$, where Class 00 is the building construction age “Before 1930”, Class $1$ the period 1930-1945, Class $2$ the epoch 1946-1960, Class $3$ the period 1961-1976, Class $4$ the time epoch 1977-1992, Class $5$ the period 1993-2006, and Class $6$ “After 2006”. In Figs. 4 and 5, we show the number of image samples in the dataset MyCD per country and building construction epoch class, as well as per city.
II-B The task of predicting the age of buildings
The multi-modal dataset MyCD was used in the challenge that we organized in 2024 to predict the age of buildings in several different cities in Europe. The goal of the challenge was to achieve the best performance possible on cities not included in the training data, i.e. on previously unseen cities. In addition, half of the test samples did not include the street-view modality. With this data challenge, we hope to advance the research on models’ out-of-distribution (OoD) generalization.
We present the evaluation results for the OoD generalization performance of the top-performing models from the challenge when: 1) both training and testing are performed with all the three modalities of the dataset MyCD, and 2) training is performed with the three modalities of the street-view and EO dataset and testing/ inference is performed using only the top-view satellite modalities, i.e. without the street-view images.
Figure 3: Left: Example street-view image of building from the dataset MyCD, and the corresponding top-view VHR image. The images are co-localized with respect to the specific building. Right: Example top-view image from our multi-modal dataset.
Figure 4: Number of images per country and building age class.
Figure 5: Number of images per city and building age class.
III Models and Methodologies
III-A Models developed to predict building construction year
In this section, we will discuss the models from the challenge that estimate the age of buildings. More specifically, we discuss the models that achieve the top/ winning results of the challenge (i.e. in Sec. V, in Figs. 11 and 12). In particular, in this section, we will discuss four different models that predict the construction epoch of buildings. To achieve very good/ top performance, they employ several different deep learning techniques, for example class label correlation modelling.
Label correlation. The team Creamble performs label correlation modelling to take advantage of the distance between the labels. For the seven building construction epoch classes, adjacent classes are correlated, while far apart classes are not. Soft labels are also used, as well as label smoothing and hyperparameter search. In addition, K-fold cross-validation, the recently proposed model EVA [1], and a model ensemble of $15$ models with soft-voting are also used. The model minimizes the cross-entropy loss with label smoothing (0.3). The flowchart is shown in Fig. 6, where the model ensemble that estimates the construction period of buildings is illustrated.
Figure 6: Flowchart of the model to predict the age of buildings by the team Creamble.
Figure 7: Flowchart diagram of the model by the team TelePIX AI, where the results are in Sec. V, in Figs. 11 and 12.
Prediction-level data fusion. The team TelePIX AI performs prediction-level data fusion (rather than data fusion at the feature level), as well as K-fold cross-validation with the city as the fold for each country. For this method, a mean confidence model ensemble is used. Moreover, as encoders, the model EVA-02 [2] is also utilized, as well as reBEN [3]. For classification, the cross-entropy loss is used. The flowchart diagram of this model is illustrated in Fig. 7. Hyper-parameter optimization using Bayesian optimization with Optuna (e.g. for the learning rate (lr), batch size, and lr decay factor) is also performed. In addition, here, the image pre-processing is: normalization, histogram equalization, random rotation and flipping, and random hide of the street-view images, where the latter is for the missing modality problem during inference (i.e. missing street-view images). For this approach, the key factors for good model performance are: powerful encoders, random hide street-view (P: 0.5) so that the model learns the significant features from the VHR images and Sentinel-2, split based on country, ensemble technique and pre-processing to increase robustness of the final prediction and training stability.
Figure 8: Flowchart of model using an additional shared feature encoder for the street- and top-view VHR images and the distribution alignment loss, by the team xmb.
Figure 9: Flowchart of the model that uses SwinT encoders, by the team The AI Buzzard.
Ensemble with country categorical feature. The team xmb trains an ensemble of four models. For this approach, a categorical feature for the country in the format of one-hot representation is used. This country categorical feature is used in the late data fusion concatenation of the features. For the three input modalities of the dataset MyCD, the latent representations after average pooling are combined, also using the country categorical feature, and then inputted to the final dense Fully Connected (FC) layer. In addition, in the ensemble, the models EfficientNet and MobileViT are also used, and geometric data fusion is also performed. The latent feature representations before average pooling are combined according to the method presented in [4], and then inputted to average pooling and, then, concatenated using the country categorical feature. To address the problem of the missing modality during inference, street-view images during training are randomly replaced with all zeros, i.e. input dropout, and shared feature modelling [5] is performed. Here, an additional shared encoder for the two different modalities of street-view and top-view VHR is used. The flowchart diagram is shown in Fig. 8. The distribution alignment loss function term [5] is also added in the objective function minimized during training.
SwinT-encoder late data fusion with attention. The model Swin Transformer (SwinT) pretrained on ImageNet22K is used by the team The AI Buzzard, for the encoders of the three different input modalities. Late data fusion using attention is performed, as well as 5-fold cross-validation training, grouped by cities and stratified by the class label. Data fusion is performed by computing a weighted sum of the embeddings. In this way, the model is able to handle missing data during inference (i.e. missing street-view images). The flowchart diagram is shown in Fig. 9. For this method, the models are first trained on all the data and, then, they are fine-tuned on the individual countries. For data augmentation, random flipping, color jittering, and random resized cropping are utilized.
Discussion about the key features. The main characteristics of the models discussed in the preceding paragraphs are: i) very good/ top pretrained model, i.e. EVA [1] or EVA-02 [2], that is used as a starting point so as to not train from scratch, ii) model ensemble, iii) hyper-parameter optimization using for example Optuna, iv) label correlation modelling between adjacent classes (per modality) because the distance between adjacent classes is smaller than the distance between classes that are not adjacent, and v) missing modality distribution alignment loss together with a shared feature encoder. Here, additional key features of the main models are: K-fold cross-validation training, geometric data fusion using the methodology in [4], and use of a categorical feature for the country in the form of one-hot vector representation.
III-B Baseline model: Age of Buildings for Sustainability (ABS)
We present the baseline model Age of Buildings for Sustainability (ABS) which is based on Transformers, and more specifically on the recently proposed model SegFormer [6, 7], which uses multi-scale features. This baseline model of the challenge employs three encoder networks, one for each of the modalities of the dataset MyCD, and performs late data fusion by concatenating the feature representations in the latent space. The flowchart of the model ABS is shown in Fig. 10. The output is the construction epoch class for the queried building. We also release our code for ABS for reproducibility [^1] The GitHub repository is: http://github.com/AI4EO/MapYourCity..
Figure 10: Flowchart of the model ABS, the baseline of the challenge we organized using the multi-modal dataset MyCD.
Figure 11: Evaluation results of the different models, and more specifically the private leaderboard of the challenge.
Figure 12: Results from the evaluation of the models, and more specifically the public leaderboard of the challenge.
IV Related Work
Many different AI models from the recent literature perform data fusion of street-view images and top-view satellite images. With AI and deep learning methods, by performing late data fusion (i.e. concatenating latent representations in the feature space), it is possible to achieve better performance than early data fusion, because street-view images and top-view satellite data have many differences, for example in resolution, object characteristics and scale, and significant differences in appearance. Street-view and top-view satellite images do not look alike and top-view data cover a larger geographical area. In general, the effective fusion of street-view ground images and nadir-view satellite or aerial images is nowadays an open research problem. By jointly using both street- and top-view images, the model proposed in [8] performs mapping of building functions using a specific data fusion strategy that combines in a decision-level manner different models trained for each image modality independently. In this way, by performing decision-level data fusion of the two modalities, the method presented in [8] combines the predictions of different models in order to accurately perform mapping of building functions, for example commercial, residential, public, and industrial, i.e. classification using four different classes.
Furthermore, nowadays, the importance of street-view images has become evident in the literature [9]. Street-view imagery is an important data source for urban analytics, helping us derive insights for making informed decisions [10]. Using street-view images, we can analyze for example the built environment, the vegetation, and the transportation [9], and the results of the analyses can be linked to important health, urban, and socio-economic studies. In this work, street-view images from Mapillary are used in the multi-modal dataset MyCD.
Street-view imagery is also important for urban infrastructure assessment [11] and urban mobility, as well as for spatial data infrastructures, urban planning [12], and urban greenery. In [11], the dataset Global Streetscapes is introduced which contains $10$ million street-view images from $688$ cities around the Earth, from $212$ countries and different regions. These images have been crowdsourced from Mapillary and KartaView. Moreover, the dataset Global Streetscapes [11] also includes metadata information for every street-view image, such as geo-location, longitude and latitude, acquisition time, and contextual, semantic, and perceptual information.
Transformer models have shown to achieve good/ top performance and outperform other models, for example models that use an architecture based on Convolutional Neural Networks (CNN), U-Net, and/or Residual Networks (ResNet). Importantly, in [13], geo-location matching of street-view images and top-view aerial images is performed using a Transformer-based model and the mechanism of attention. The model TransGeo proposed in [13] is based on Transformers because global information modeling is crucial. It performs attention-guided non-uniform cropping of images as a means of attending and zooming-in on only informative image patches, not focusing on non-informative (for the task) parts of the image.
Different datasets that contain top-view satellite or aerial images exist, for example the dataset OpenEarthMap [14]. This publicly available dataset is designed for the task of high-resolution land cover mapping semantic segmentation. More specifically, OpenEarthMap contains $5000$ aerial and satellite images, and for each image, it has land cover labels/ annotations for $8$ different classes. It covers in total $97$ regions, from $44$ countries across $6$ continents. This recent dataset OpenEarthMap has been used, for example, in the model proposed in [15] which performs semantic segmentation land cover classification in a few-shot continual learning setting.
V Evaluation and Results
Evaluation metrics. In this work, for the numerical evaluation results, we compute the classification confusion matrix for the task of predicting the construction period of buildings, where we have $7$ different classes, for example “After 2006”. We calculate the average of the diagonal items of the confusion matrix, and we report this metric, which is also known as the Mean Producer Accuracy (MPA) evaluation metric.
We present the evaluation of the models from the challenge in Figs. 11 and 12. We use MPA_ALL_MODALITIES to denote the setting when both training and testing are performed with all modalities. In addition, we also use MPA_TOP_VIEW to denote the setting when training is performed with all modalities and inference is performed using only the two top-view modalities. TOTAL is the average of the two MPAs.
In Figs. 11 and 12, in the results [^2] https://platform.ai4eo.eu/map-your-city/leaderboard, the MPA_TOP_VIEW scores are lower than MPA_ALL_MODALITIES. The top score for MPA_ALL_MODALITIES in the private leaderboard is $0.74$ indicating that the estimation of the construction epoch of buildings in unseen cities is possible. When using only the top-view data during inference, then the performance of the models decreases by more than $10$ points in the MPA score. Finally, we observe that for the TOTAL scores, the top winning values are approximately $0.65$ in the private leaderboard.
VI Conclusion
In this work, we have introduced the new multi-modal benchmark dataset MyCD and the real-world multi-modal application study for EO generalization. The performance evaluation of the models from the challenge show that estimating the age of buildings in previously unseen cities is possible and effective, as well as predicting the construction period of buildings/ houses from only top-view satellite images, i.e. without using the street-view images during inference.
Acknowledgment
The challenge was accomplished thanks to the contributions of many people. We want to thank all the team: Juan Pedro, Dennis Albrecht, and everyone in ESA AI4EO.

### v2

@@
 In this paper, we introduce a new benchmark multi-modal dataset, i.e.
-the Map your City Dataset (MyCD), containing top-view Very High Resolution (VHR) images, Earth Observation (EO) multi-spectral data from the Copernicus Sentinel-2 satellite constellation, and street-view images in many different cities in Europe, co-localized with respect to the building under study and labelled with the construction epoch.
+the Map your City Dataset (MyCD), containing top-view Very High Resolution (VHR) images, Earth Observation (EO) multi-spectral data from the Copernicus Sentinel-2 satellite constellation, and street-view images in many different cities in Europe that are co-localized with respect to the building under study and labelled with the construction epoch.
 We assess EO generalization performance on new/ previously unseen cities that have been held-out from training and appear only during inference.
 In this work, we present the community-based data challenge we organized based on MyCD.
-The AI4EO Challenge MapYourCity was opened in 2024 for 4 months.
-We present the Top-4 performing models, and the evaluation results.
-During inference, the performance of the models using both all three input modalities and only the two top-view modalities, i.e.
-without the street-view images, is examined.
-The evaluation results show that the models are effective and can achieve good performance on this difficult real-world task of estimating the age of buildings, even on previously unseen cities, as well as even using only the two top-view modalities (i.e.
+The AI4EO Challenge ESA MapYourCity was opened in 2024 for 4 months.
+In this paper, we present the Top-4 performing models of the challenge, and the evaluation results.
+During inference, the performance of the models using: i) both all three input modalities, and ii) only the two top-view modalities, i.e.
+without the street-view ground images, is examined.
+The evaluation results in this work show that the models to estimate the construction year of buildings are effective and can achieve good performance on this difficult important real-world task, even when inference is on previously unseen cities, as well as even when using only the two top-view modalities (i.e.
 VHR and Sentinel-2) during inference.
@@
 Sustainable buildings minimize energy consumption and are a key part of responsible and sustainable urban planning and development that seeks to effectively combat climate change, as well as rising temperatures.
-With increase in urbanization and Earth Sciences research into urban areas, the need to estimate the energy efficiency of buildings, so as to meet important sustainability goals, has never been greater.
+With an increase in urbanization and Earth sciences research in urban areas, the need to estimate the energy efficiency of buildings, to meet important sustainability goals, has never been greater.
 The age of buildings plays an important role in energy modelling, as well as in urban policies.
-The age of the building can be used as a proxy for the energy efficiency.
+The age of buildings can be used as a proxy for energy efficiency.
 However, data for the construction period of buildings is not currently available globally in a consistent and homogeneous manner.
-To this end, as a start, in this study, we focus on European cities.
-In this real-world application study for sustainability and climate, the main aim is to develop models that automatically estimate the construction year of buildings with improved performance using Artificial Intelligence (AI) techniques and cross-view datasets based on street-view images and Earth Observation (EO) datasets.
-The goals of this paper are to present the new multi-modal dataset, examine example images from this street-view and EO dataset, present the community data challenge that we organized, and present the evaluation results of the top-performing models from the challenge.
+To this end, as a starting point, in this paper, we focus on European cities.
+In this real-world application study for sustainability and climate, the main objective is to develop models that automatically estimate the construction year of buildings with improved performance using Artificial Intelligence (AI) techniques and cross-view datasets based on street-view images and Earth Observation (EO) datasets.
+The main goals of this paper are to present the new multi-modal dataset, examine example images from this street-view and EO dataset, present the community data challenge that we organized, and present the evaluation results of the four top-performing models from the challenge.
 We introduce a new benchmark dataset for a real-world multi-modal application study.
-It consists of three modalities, including street-view ground images, top-view satellite Very High Resolution (VHR) images at 50 cm resolution, and Copernicus Sentinel-2 multi-spectral satellite data.
+It consists of three modalities, including street-view ground images, top-view satellite Very High Resolution (VHR) images at 50 cm resolution, and Copernicus Sentinel-2 multi-spectral satellite data, i.e.
+all spectral bands.
 All the modalities are co-localized with respect to the specific building under study, as well as its label of the construction epoch.
-The problem is formulated as a classification task with seven classes for the construction epoch, ranging from 1900 to nowadays.
+The novel multi-modal benchmark dataset is suitable for correlation learning between the different modalities, as well as for data fusion.
+The problem is formulated as a classification task with seven classes for the construction epoch of buildings, ranging from 1900 to nowadays.
 This paper evaluates the EO generalization performance of the models on cities that are not included in the training dataset, i.e.
@@
 We examine the performance of the models when both training and testing are performed with all 3 modalities.
-There is benefit in using the top-view VHR images (and to a lesser extent, Sentinel-2 data), in addition to the street-view images.
-In this work, we also examine the performance of the models when training is performed with all the three modalities, and testing is performed without the street-view images, i.e.
+There is benefit in using the top-view VHR images (and, to a lesser extent, Sentinel-2 data), in addition to the street-view images.
+In this work, we also examine the performance of the models when training is performed with all three modalities, and testing is performed without the street-view images, i.e.
 when using only two out of the three encoders (i.e.
 top-view VHR and Sentinel-2 images).
-These results are compared to those when training and testing are performed using all the three modalities.
+In this paper, we compare these results to those when training and testing are both performed using all three modalities.
 Figure 1: Left: Example street-view image from the multi-modal dataset. Right: Example top-view satellite VHR image.
-Main contributions.
-Our contributions are that this work introduces a novel multi-modal dataset, presents the challenge we organized to estimate the age of buildings in different cities in Europe using this dataset, examines the performance evaluation results of the models from the challenge, and shows the potential of effectively combining and fusing street-view, top-view satellite and remotely sensed Sentinel-2 EO data for accurately predicting the construction period of buildings.
-In this paper, we present and compare the performance of the models from the challenge that have been trained and tested on the new multi-modal street-view and EO dataset.
-II Data and Task
-II-A The multi-modal street-view and EO dataset MyCD
-We present the multi-modal dataset and examine example images from this street-view and EO dataset.
-The dataset MyCD encompasses the three modalities of: a) Street-view images, b) Top-view satellite VHR RGB images at 50 cm resolution, and c) Multi-spectral 12-band S-2 data.
-These input modalities have very different characteristics (number of spectral bands, spatial resolution, and building perspectives from top and street-view).
-In the dataset, the three modalities are co-localized with respect to the specific building under study.
-In Fig.
-1, we show two example images for the modalities of street-view ground and top-view satellite VHR, respectively.
 Figure 2: Example top-view satellite VHR image and the corresponding geo-localized S-2 image (here, RGB bands for visualization purposes, where the dataset includes all the $12$ spectral bands, i.e.
 L2A) from the multi-modal dataset MyCD.
-In Fig.
-2, an example top-view VHR image and the corresponding S-2 image, both from the new multi-modal dataset MyCD, are shown, together with the building location in the centre, as a blue dot in the image.
-We observe the geospatial overlay of the two top-view image modalities and that the two top-view modalities are centered at the building under study.
-In Fig. 3, additional examples of street-view and top-view images from the new multi-modal dataset MyCD are shown.
-The dataset MyCD comprises cross-view image data from $6$ countries in Europe, and more specifically from $19$ cities, for $7$ different building construction epoch classes extracted from the EUBUCCO database.
-For the training set data, $15$ cities have been chosen, while for the test set data, the remaining $4$ cities have been used.
-The building construction epoch classes are $7$, i.e.
-labelled from 00 to $6$, where Class 00 is the building construction age “Before 1930”, Class $1$ the period 1930-1945, Class $2$ the epoch 1946-1960, Class $3$ the period 1961-1976, Class $4$ the time epoch 1977-1992, Class $5$ the period 1993-2006, and Class $6$ “After 2006”.
-In Figs.
-4 and 5, we show the number of image samples in the dataset MyCD per country and building construction epoch class, as well as per city.
-II-B The task of predicting the age of buildings
-The multi-modal dataset MyCD was used in the challenge that we organized in 2024 to predict the age of buildings in several different cities in Europe.
-The goal of the challenge was to achieve the best performance possible on cities not included in the training data, i.e.
-on previously unseen cities.
-In addition, half of the test samples did not include the street-view modality.
-With this data challenge, we hope to advance the research on models’ out-of-distribution (OoD) generalization.
-We present the evaluation results for the OoD generalization performance of the top-performing models from the challenge when: 1) both training and testing are performed with all the three modalities of the dataset MyCD, and 2) training is performed with the three modalities of the street-view and EO dataset and testing/ inference is performed using only the top-view satellite modalities, i.e.
-without the street-view images.
-Figure 3: Left: Example street-view image of building from the dataset MyCD, and the corresponding top-view VHR image.
-The images are co-localized with respect to the specific building.
-Right: Example top-view image from our multi-modal dataset.
-Figure 4: Number of images per country and building age class.
-Figure 5: Number of images per city and building age class.
-III Models and Methodologies
-III-A Models developed to predict building construction year
-In this section, we will discuss the models from the challenge that estimate the age of buildings.
-More specifically, we discuss the models that achieve the top/ winning results of the challenge (i.e.
-in Sec.
-V, in Figs.
-11 and 12).
-In particular, in this section, we will discuss four different models that predict the construction epoch of buildings.
-To achieve very good/ top performance, they employ several different deep learning techniques, for example class label correlation modelling.
-Label correlation.
-The team Creamble performs label correlation modelling to take advantage of the distance between the labels.
-For the seven building construction epoch classes, adjacent classes are correlated, while far apart classes are not.
-Soft labels are also used, as well as label smoothing and hyperparameter search.
-In addition, K-fold cross-validation, the recently proposed model EVA [1], and a model ensemble of $15$ models with soft-voting are also used.
-The model minimizes the cross-entropy loss with label smoothing (0.3).
-The flowchart is shown in Fig.
-6, where the model ensemble that estimates the construction period of buildings is illustrated.
-Figure 6: Flowchart of the model to predict the age of buildings by the team Creamble.
-Figure 7: Flowchart diagram of the model by the team TelePIX AI, where the results are in Sec. V, in Figs. 11 and 12.
-Prediction-level data fusion.
-The team TelePIX AI performs prediction-level data fusion (rather than data fusion at the feature level), as well as K-fold cross-validation with the city as the fold for each country.
-For this method, a mean confidence model ensemble is used.
-Moreover, as encoders, the model EVA-02 [2] is also utilized, as well as reBEN [3].
-For classification, the cross-entropy loss is used.
-The flowchart diagram of this model is illustrated in Fig.
-7.
-Hyper-parameter optimization using Bayesian optimization with Optuna (e.g.
-for the learning rate (lr), batch size, and lr decay factor) is also performed.
-In addition, here, the image pre-processing is: normalization, histogram equalization, random rotation and flipping, and random hide of the street-view images, where the latter is for the missing modality problem during inference (i.e.
-missing street-view images).
-For this approach, the key factors for good model performance are: powerful encoders, random hide street-view (P: 0.5) so that the model learns the significant features from the VHR images and Sentinel-2, split based on country, ensemble technique and pre-processing to increase robustness of the final prediction and training stability.
-Figure 8: Flowchart of model using an additional shared feature encoder for the street- and top-view VHR images and the distribution alignment loss, by the team xmb.
-Figure 9: Flowchart of the model that uses SwinT encoders, by the team The AI Buzzard.
-Ensemble with country categorical feature.
-The team xmb trains an ensemble of four models.
-For this approach, a categorical feature for the country in the format of one-hot representation is used.
-This country categorical feature is used in the late data fusion concatenation of the features.
-For the three input modalities of the dataset MyCD, the latent representations after average pooling are combined, also using the country categorical feature, and then inputted to the final dense Fully Connected (FC) layer.
-In addition, in the ensemble, the models EfficientNet and MobileViT are also used, and geometric data fusion is also performed.
-The latent feature representations before average pooling are combined according to the method presented in [4], and then inputted to average pooling and, then, concatenated using the country categorical feature.
-To address the problem of the missing modality during inference, street-view images during training are randomly replaced with all zeros, i.e.
-input dropout, and shared feature modelling [5] is performed.
-Here, an additional shared encoder for the two different modalities of street-view and top-view VHR is used.
-The flowchart diagram is shown in Fig.
-8.
-The distribution alignment loss function term [5] is also added in the objective function minimized during training.
-SwinT-encoder late data fusion with attention.
-The model Swin Transformer (SwinT) pretrained on ImageNet22K is used by the team The AI Buzzard, for the encoders of the three different input modalities.
-Late data fusion using attention is performed, as well as 5-fold cross-validation training, grouped by cities and stratified by the class label.
-Data fusion is performed by computing a weighted sum of the embeddings.
-In this way, the model is able to handle missing data during inference (i.e.
-missing street-view images).
-The flowchart diagram is shown in Fig.
-9.
-For this method, the models are first trained on all the data and, then, they are fine-tuned on the individual countries.
-For data augmentation, random flipping, color jittering, and random resized cropping are utilized.
-Discussion about the key features.
-The main characteristics of the models discussed in the preceding paragraphs are: i) very good/ top pretrained model, i.e.
-EVA [1] or EVA-02 [2], that is used as a starting point so as to not train from scratch, ii) model ensemble, iii) hyper-parameter optimization using for example Optuna, iv) label correlation modelling between adjacent classes (per modality) because the distance between adjacent classes is smaller than the distance between classes that are not adjacent, and v) missing modality distribution alignment loss together with a shared feature encoder.
-Here, additional key features of the main models are: K-fold cross-validation training, geometric data fusion using the methodology in [4], and use of a categorical feature for the country in the form of one-hot vector representation.
-III-B Baseline model: Age of Buildings for Sustainability (ABS)
-We present the baseline model Age of Buildings for Sustainability (ABS) which is based on Transformers, and more specifically on the recently proposed model SegFormer [6, 7], which uses multi-scale features.
-This baseline model of the challenge employs three encoder networks, one for each of the modalities of the dataset MyCD, and performs late data fusion by concatenating the feature representations in the latent space.
-The flowchart of the model ABS is shown in Fig.
-10.
-The output is the construction epoch class for the queried building.
-We also release our code for ABS for reproducibility [^1] The GitHub repository is: http://github.com/AI4EO/MapYourCity..
-Figure 10: Flowchart of the model ABS, the baseline of the challenge we organized using the multi-modal dataset MyCD.
-Figure 11: Evaluation results of the different models, and more specifically the private leaderboard of the challenge.
-Figure 12: Results from the evaluation of the models, and more specifically the public leaderboard of the challenge.
-IV Related Work
+Main contributions.
+Our contributions are that this work introduces a novel multi-modal dataset, presents the challenge we organized to estimate the age of buildings in different cities in Europe using this dataset, examines the performance evaluation results of the four top models from the challenge, and shows the potential of effectively combining and fusing street-view, top-view satellite and remotely sensed Sentinel-2 EO data for accurately predicting the construction period of buildings which is crucial for sustainability, energy efficiency, urban planning and development, and safety.
+In this paper, we present and compare the performance of the models from the challenge that have been trained and evaluated on the new multi-modal street-view and EO top-view labelled dataset.
+II Related Work
 Many different AI models from the recent literature perform data fusion of street-view images and top-view satellite images.
@@
 concatenating latent representations in the feature space), it is possible to achieve better performance than early data fusion, because street-view images and top-view satellite data have many differences, for example in resolution, object characteristics and scale, and significant differences in appearance.
-Street-view and top-view satellite images do not look alike and top-view data cover a larger geographical area.
+Street-view and top-view satellite images do not look alike, and top-view data cover a larger geographical area.
 In general, the effective fusion of street-view ground images and nadir-view satellite or aerial images is nowadays an open research problem.
-By jointly using both street- and top-view images, the model proposed in [8] performs mapping of building functions using a specific data fusion strategy that combines in a decision-level manner different models trained for each image modality independently.
+By jointly using both street-view and top-view images, the model proposed in [1] performs mapping of building functions using a specific data fusion strategy that combines in a decision-level manner different models trained for each image modality independently.
 In this way, by performing decision-level data fusion of the two modalities, the method presented in [1] combines the predictions of different models in order to accurately perform mapping of building functions, for example commercial, residential, public, and industrial, i.e.
@@
 Furthermore, nowadays, the importance of street-view images has become evident in the literature [2].
-Street-view imagery is an important data source for urban analytics, helping us derive insights for making informed decisions [10].
-Using street-view images, we can analyze for example the built environment, the vegetation, and the transportation [9], and the results of the analyses can be linked to important health, urban, and socio-economic studies.
+Street-view imagery is an important data source for urban analytics, helping us to make informed decisions [3].
+Using street-view images, we can analyze for example the built environment, vegetation, and transportation [2], and the results of these analyses can be linked to important health, urban, socio-economic, development, and sustainability studies.
 In this work, street-view images from Mapillary are used in the multi-modal dataset MyCD.
@@
 The model TransGeo proposed in [6] is based on Transformers because global information modeling is crucial.
-It performs attention-guided non-uniform cropping of images as a means of attending and zooming-in on only informative image patches, not focusing on non-informative (for the task) parts of the image.
+It performs attention-guided non-uniform cropping of images as a means of attending and zooming-in on only informative image patches, not focusing on noninformative (for the task) parts of the image.
+Figure 3: Left: Example street-view image of building from the dataset MyCD, and the corresponding top-view VHR image.
+The images are co-localized with respect to the specific building.
+Right: Example top-view image from our multi-modal dataset.
+Figure 4: Number of images per country and building age class.
+Figure 5: Number of images per city and building age class.
 Different datasets that contain top-view satellite or aerial images exist, for example the dataset OpenEarthMap [7].
@@
 It covers in total $97$ regions, from $44$ countries across $6$ continents.
-This recent dataset OpenEarthMap has been used, for example, in the model proposed in [15] which performs semantic segmentation land cover classification in a few-shot continual learning setting.
-V Evaluation and Results
-Evaluation metrics.
-In this work, for the numerical evaluation results, we compute the classification confusion matrix for the task of predicting the construction period of buildings, where we have $7$ different classes, for example “After 2006”.
+This recent dataset OpenEarthMap has been used, for example, in the model proposed in [8] that performs semantic segmentation land cover classification in a few-shot continual learning setting.
+III Data and Task
+III-A The multi-modal street-view and EO dataset MyCD
+In this work, we present the multi-modal dataset and examine example images from this street-view and EO dataset.
+The dataset MyCD encompasses the three modalities of: a) Street-view images, b) Top-view satellite VHR RGB images at 50 cm resolution, and c) Multi-spectral 12-band S-2 data.
+These input modalities have very different characteristics (number of spectral bands, spatial resolution, and building perspectives from top and street-view).
+In the dataset, the three modalities are co-localized with respect to the specific building under study.
+In Fig.
+1, we show two example images for the modalities of street-view ground and top-view satellite VHR, respectively.
+In Fig.
+2, an example top-view VHR image and the corresponding S-2 image, both from the new multi-modal dataset MyCD, are shown, together with the building location in the centre, as a blue dot in the image.
+We observe the geospatial overlay of the two top-view image modalities and that the two top-view modalities are centered at the building under study.
+In Fig.
+3, additional examples of street-view and top-view images from the new multi-modal dataset MyCD are shown.
+The dataset MyCD comprises cross-view image data from $6$ countries in Europe, and more specifically from $19$ cities, for $7$ different building construction epoch classes extracted from the EUBUCCO database.
+For the training set data, $15$ cities have been chosen, while for the test set data, the remaining $4$ cities have been used.
+The building construction epoch classes are $7$, i.e.
+labelled from 00 to $6$, where Class 00 is the building construction age “Before 1930”, Class $1$ the period 1930-1945, Class $2$ the epoch 1946-1960, Class $3$ the period 1961-1976, Class $4$ the time epoch 1977-1992, Class $5$ the period 1993-2006, and Class $6$ “After 2006”.
+In Figs.
+4 and 5, we show the number of image samples in the dataset MyCD per country and building construction epoch class, as well as per city.
+For the new multi-modal dataset and for the challenge “MapYourCity”, the training dataset includes data from $15$ cities in Europe, while the evaluation dataset (i.e.
+Private Leaderboard and Public Leaderboard) contains images from $4$ additional cities.
+The Private and Public Leaderboards include data in a random split from the above mentioned $4$ cities.
+Figure 6: Evaluation results of the different models, and more specifically the private leaderboard of the challenge.
+Figure 7: Results from the evaluation of the models, and more specifically the public leaderboard of the challenge.
+III-B The task of predicting the age of buildings
+The multi-modal dataset MyCD is used in the challenge that we organized in 2024 to predict the age of buildings in several different cities in Europe.
+The goal of the challenge is to achieve the best performance possible on cities not included in the training data, i.e.
+on previously unseen cities.
+In addition, half of the test samples did not include the street-view modality.
+With this data challenge, we hope to advance the research on models’ out-of-distribution (OoD) generalization.
+We present the evaluation results for the OoD generalization performance of the top-performing models from the challenge when: 1) both training and testing are performed with all the three modalities of the dataset MyCD, and 2) training is performed with the three modalities of the street-view and EO dataset and testing/ inference is performed using only the top-view satellite modalities, i.e.
+without the street-view images.
+IV Models developed to predict the construction year of buildings
+In the following sections, we discuss models from the challenge that estimate the age of buildings.
+More specifically, we discuss the four models that achieve the top/ winning results of the challenge (in Figs.
+6 and 7).
+To achieve very good/ top performance, the models employ several different deep learning techniques, for example, class label correlation modelling, choosing the best performing model architectures, using a different encoder for feature extraction for each modality or a shared encoder to align and share features from different modalities, or training different models for different countries.
+Moreover, the top models perform ensemble techniques, weight the importance of features, employ K-fold Cross Validation, perform late fusion to combine the extracted features, combine different feature fusion methods to learn feature importance, i.e.
+feature concatenation and/or geometric data fusion, ensemble the outputs of different trained models across different folds, or employ a hide-out strategy for the street-view images to ensure model robustness.
+The evaluation metric used is the Mean Percentage Accuracy (MPA).
+For each model, we present the methodology, data preparation, models used, ensemble approach, experiments, results and discussion.
+Also, we present the baseline model Age of Buildings for Sustainability (ABS), which is based on recent Transformer models.
+V The first top winner team: Creamble
+This section presents the first top solution to the AI4EO MapYourCity Challenge.
+The task involves classifying building construction years using three input images: street view, orthophoto, and Sentinel-2.
+V-A Methodology
+In this section, we present the approach of the team that wins the AI Challenge “MapYourCity”, which focuses on predicting building ages from three distinct domain images: Street-View, Top-View, and Sentinel-2.
+The methodology involves carefully handling label correlation by converting labels into soft labels based on Gaussian Probabilistic Distribution, guided by KL-Divergence Loss.
+This approach leads to initial improvements, which were further enhanced using Label Smoothing, with a hyperparameter optimization of $0.3$ yielding the best results.
+Figure 8: Multi-Modal and Folds Ensemble with Soft-Voting Approach: Flowchart of the model to predict the age of buildings by the team Creamble, i.e. the first top winner team.
+Figure 9: Label Correlation.
+For image preparation, size distributions are analyzed and aspect ratios are selected that minimize the loss during resizing.
+Experimenting with various different state-of-the-art models, the Transformer-based EVA-02 model is chosen for its superior performance in feature extraction across diverse image domains.
+Ensemble techniques are integral to this success; a soft-voting approach is utilized, combining predictions across models and folds.
+Additionally, ensemble layers are trained by freezing domain-specific models and focusing on ensembling logits, which significantly improves the final decision-making process.
+The experimental design incorporates K-fold Cross Validation, with model selection guided by Mean Percentage Accuracy (MPA).
+Optimization techniques such as CosineAnnealingWarmRestarts and the use of AdamW help prevent overfitting, ensuring robust performance.
+The flowchart is shown in Fig.
+8, where the model ensemble that estimates the construction period of buildings is illustrated.
+Figure 10: Street-View Best Model Training Logs.
+V-B Data Preparation
+Label Correlation.
+The target of this challenge is to predict the year a building was constructed, which means that there is a correlation between labels based on temporal proximity.
+For instance, Class 6 “after 2006” is closer to Class 5 “1993-2006” than Class 0 “before 1930.” Therefore, considering this correlation is crucial in the learning process.
+Our first approach is to assume that the label distribution is probabilistic, rather than binary.
+We convert the labels into soft labels using a Gaussian probabilistic distribution, with the learning guided by the KL-Divergence Loss, in Fig.
+9.
+This approach results in a slight improvement.
+However, because of the manual adjustment of the variance parameter, some uncertainty and cost are introduced.
+To overcome this, we implement Label Smoothing and achieved our best performance with a smoothing factor of $0.3$, which is determined through hyperparameter search.
+Image Size Distribution.
+For resizing street-view images, we first analyze the image size distribution.
+We select a 2:3 aspect ratio as it best represents the overall distribution, allowing for resizing with minimal loss.
+For top-view and Sentinel-2 images, which have a 1:1 aspect ratio, we follow the training size of the model during resizing.
+V-C Models
+We evaluate several high-performing models on ImageNet for image classification tasks.
+Specifically, we test Transformer-based models such as EVA [9, 10], BEiT [11], and MaxViT [12], which perform well.
+However, performance varies between different image types (street-view, top-view, and Sentinel-2), likely due to the datasets and pretraining methods used.
+Ultimately, we select EVA-02 [10] as the best model for this task, as it achieves superior performance in feature extraction for this multi-modal dataset, given its parameter size and the nature of the images.
+V-D Ensemble
+We apply a soft-voting ensemble approach to accommodate the test dataset’s structure, which has two cases: availability of all data sources and absence of street-view images.
+The method, depicted in Figure 9, efficiently handles both cases and incorporates K-fold Cross Validation.
+This avoids the need for additional layers to combine multi-modal data, improving efficiency and reducing overfitting on the public leaderboard.
+V-E Results and Discussion
+Experimental Design.
+We utilize K-fold Cross Validation and integrate it into the ensemble algorithm.
+The competition’s Mean Percentage Accuracy (MPA), which closely resembles recall, is used as the criterion to select the best model.
+Figure 11: Overall diagram of our approach: training, inference and ensemble: Flowchart diagram of the model by the team TelePIX AI, i.e.
+the second top winner team, where the results are in Figs.
+6 and 7.
+Figure 12: Ten sample orthophoto image from each city on the same country (separated by blue line) from the top: PNN, HUN, FMW, and QCD.
+Detailed Results and Discussion.
+During optimization, we observe that the highest validation recall values appeared at two points: early in epoch 4 and late in epoch 7.
+The difference between these two points was due to the gap between training and validation recall.
+For example, the validation scores at epochs 4 and 7 were 0.7297 and 0.7290, respectively, while the training recall values were 0.92 and 0.96.
+This suggested potential overfitting in later epochs.
+However, this pattern was inconsistent across folds (Fig.
+10).
+To mitigate this, we employ CosineAnnealingWarmRestarts, allowing us to adjust the learning rate regularly, preventing local minima and overfitting.
+AdamW was also used to improve generalization.
+Among the ensemble methods, the combination of the results of regression and classification using Cross-Entropy is also successful.
+The Euclidean distance from the MSE model predictions is converted into a percentage and incorporated into the Cross-Entropy Softmax values, boosting performance.
+However, this ensemble structure was unfortunately not included in the final submission.
+Nonetheless, there is potential for its application in problems that can be interpreted from both regression and classification perspectives, where labels have some degree of correlation.
+VI The second top winner team: TelePIX AI
+This section presents the second top solution to the AI4EO MapYourCity Challenge.
+The task involves classifying building construction years using three input images: street-view, orthophoto, and Sentinel-2.
+Prediction-level data fusion.
+The team TelePIX AI performs prediction-level data fusion (rather than data fusion at the feature level), as well as K-fold cross-validation with the city as the fold for each country.
+For this method, a mean confidence model ensemble is used.
+Moreover, as encoders, the model EVA-02 [10] is also utilized, as well as reBEN [13].
+For classification, the cross-entropy loss is used.
+The flowchart diagram of this model is illustrated in Fig.
+11.
+Hyper-parameter optimization using Bayesian optimization with Optuna (e.g., for the learning rate (lr), batch size, and the lr decay factor) is also performed.
+In addition, here, the image pre-processing block is: normalization, histogram equalization, random rotation and flipping, and random hide of the street-view ground images, where the latter is for the missing modality problem during inference (i.e.
+missing street-viewimages).
+For this approach, the key factors for good model performance are: powerful encoders, random hide street-view (P: 0.5) so that the model learns the significant features from the VHR images and Sentinel-2, split based on country, ensemble technique and pre-processing to increase robustness of the final prediction and training stability.
+VI-A Methodology
+To address the challenge MapYourCity, this approach involves careful data preparation and model design.
+The proposed model is trained separately for each country to reduce task complexity, using City-Fold cross-validation for data splitting.
+The model, referred to as TBBE2F, processes each input image through pre-trained encoders that are fine-tuned during training.
+Various encoders are explored, and the EVA02 models are selected as the best performing for street-view and orthophoto encoder and reBEN as encoder for Sentinel-2 inputs.
+A hide-out strategy is employed for the street-view ground images to ensure model robustness, even when this input is unavailable.
+Figure 13: X1 and X2 encoder performance comparison.
+Figure 14: Flowchart of model using an additional shared feature encoder for the street- and top-view VHR images and the distribution alignment loss, by the team xmb, i.e. the second best model.
+The final submission to the challenge is generated by ensembling the outputs of 15 trained models in different folds, i.e.
+City-Fold cross-validation.
+Experimental results show that this approach achieves MPA 73.89% on the private test set for all input modalities, and MPA 57.97% when using only the top-view input.
+These results demonstrate the effectiveness of the proposed encoder selection, hyperparameter tuning, and ensembly strategy.
+VI-B Data Preparation
+Data preparation aims to reduce the complexity of the task of the challenge.
+Orthophoto top-view images from cities within the same country tend to appear similar, as it can be seen in Fig.
+12, while images from different countries show significant variation.
+To take advantage of this characteristic, the model is trained separately for each country to minimize the complexity of the patterns (i.e.
+discriminative features) that are needed to be learned.
+The data are split for training and validation using K-Fold cross-validation, with each fold representing a city within the respective country.
+For the countries in the test set —FMW, PNN, HUN, and QCD— 6, 3, 3, and 3 folds are created, respectively, with the number of folds varying based on the number of cities in each country (FMW had the most).
+For each fold, 2-3 cities are used for training, while 1 city is reserved for validation.
+Several preprocessing techniques are applied, including normalization for all three input images, histogram equalization for orthophoto images, and augmentations such as random rotation and flipping.
+VI-C Models
+The architecture of the proposed model, which is referred as TBBE2F, is shown in Figure 11.
+The model processes three input images: street-view (X1), orthophoto top-view (X2), and Sentinel-2 (X3).
+Each input is handled by its own dedicated pre-trained encoder.
+Given the distinct domains of these images, separate encoders are assigned to each modality.
+During training, these encoders are fine-tuned to enhance feature extraction.
+The extracted features are represented by vectors V1, V2, and V3 for X1, X2, and X3, respectively, with the size of each vector dependent on the chosen encoder.
+A variety of encoders are explored, including SeResNext [14], MobileNet [15], EfficientNet [16], and several ViT-based [17] encoders pre-trained variations, along with EVA02 [10] and reBEN [18], which are combined across the three inputs.
+After feature extraction, the resulting vectors are passed through a fully-connected (dense) layer, producing seven logits (i.e.
+corresponding to the number of classes).
+These logits are concatenated and passed through another fully connected layer, reducing the 21 logits back to seven.
+The final predictions are processed through a SoftMax function to obtain the class probabilities.
+In addition, a hide-out strategy is applied during training for the X1 input.
+With a probability of 0.5, the X1 image is blacked out (multiplied by 0), encouraging the model to avoid over-reliance on X1 and maintain the performance when X1 was absent.
+This approach also allows the model to focus on learning important features from X2 and X3, while using X1 as a guide.
+The reason for this design is because during inference, there are instances where X1 is not available.
+In that case, a fully black image is used as a replacement for X1.
+VI-D Ensemble
+The training process results in 15 models, each corresponding to a fold from one of the four countries as described in the data preparation section.
+Inference is conducted using these models on the test datasets relevant to the country in which each model is trained on.
+The outputs for each country are averaged, and the final results are merged into a single submission file, as also depicted in Figure 11.
+VI-E Results and Discussion
+Experimental Design.
+To determine the best solution, the focus is on first finding the most suitable encoder for each input type.
+This is done by varying the encoder for X1 while keeping the encoders for X2 and X3 fixed, then varying the encoder for X2 while keeping X1 and X3 fixed, and similarly for X3 while keeping X1 and X2 fixed.
+The model is trained using the training and validation data split, as described in the data preparation section.
+Once the best combination of encoders is identified, Optuna is employed to perform hyperparameter search.
+The variables tuned during this search include batch size, learning rate, warm-up steps, and decay factor.
+The optimal combination is then used for the inference in the test set, and the results are ensembled, as outlined in the ensemble section, to generate the final submission.
+Detailed Results and Discussion.
+The experimental results show that the encoder that performs well for X1 also performs well for X2, and this is important.
+A comparison of the performance of each encoder type can be seen in Figure 13.
+The EVA02 encoder emerges as the best, with a significant margin over the other architectures.
+Given that larger models usually tend to yield better performance, we select the EVA02-L variant for both X1 and X2.
+For X3, because of its different number of channels, the best-performing encoder is reBEN, which utilizes all available channels.
+The optimal hyperparameters, identified through the Optuna search, are as follows: a learning rate of 1.32, a batch size of 8, 300 warm-up steps, and a decay factor of 1.1.
+The experiments also reveal that random hiding of street-view images improve the overall MPA by 2-3%, ensembling boosted it by 1-3%, and pre-processing with histogram equalization contribute an additional 1-2% improvement.
+Ultimately, the proposed solution achieves MPA 73.89% on the private leaderboard/ private test set of the challenge for all modalities and MPA 57.97% for the top-view-only setting.
+VII The second best model: Team xmb
+We present the second place solution for the AI4EO MapYourCity Challenge.
+Ensemble with country categorical feature.
+The team xmb trains an ensemble of four models.
+For this approach, a categorical feature for the country in the format of one-hot representation is used.
+This country categorical feature is used in the late data fusion concatenation of the features.
+For the three input modalities of the dataset MyCD, the latent representations after average pooling are combined, also using the country categorical feature, and then inputted into the final dense fully-connected layer.
+In addition, in the ensemble, the models EfficientNet and MobileViT are also used, and geometric data fusion is also performed.
+The latent feature representations before average pooling are combined according to the method presented in [19], and then inputted to average pooling and, then, concatenated using the country categorical feature.
+To address the problem of the missing modality during inference, street-view images during training are randomly replaced with all zeros, i.e.
+input dropout, and shared feature modelling [20] is performed.
+Here, an additional shared encoder for the two different modalities of street-view and top-view VHR is used.
+The flowchart diagram is shown in Fig.
+14.
+The distribution alignment loss function term [20] is also added to the objective function that is minimized during training.
+VII-A Methodology
+The proposed approach involves two types of classification models: Type I, where both training and inference are performed using the full set of modalities, and Type II, where training is carried out with the full modalities, but inference is limited to only the two top-view modalities.
+The submission, which is named xmb, achieves an overall score of $0.6468$, which is here calculated as the average of MPA_ALL_MODALITIES and MPA_TOP_VIEW.
+The code for the solution is available in Github: http://github.com/xmba15/ai4eo_map_your_city.
+VII-B Data Preparation
+As shown in Figure 15, there is a slight class imbalance in the training dataset, with label 0 having twice as many samples as the other labels.
+We address this imbalance by using FocalLoss during training for all of our models.
+Figure 15: Distribution of the Training Dataset Labels.
+Both the training and test sets include information about six countries.
+Assuming that each country has distinct styles of houses, which can help us understand the age of buildings in different regions, we convert the country index into a one-hot encoded vector and use it as a categorical feature.
+We do not apply this approach to cities because the test dataset can have a held-out city.
+VII-C Model(s)
+We propose here two types of classification models.
+For type I, both training and inference are carried out using all available modalities.
+Three inputs from these modalities are processed through three encoders of the same feature extraction model and then combined by late fusion.
+Additionally, we employ two data fusion methods: Feature Concatenation (Figure 16) and Geometric Fusion (Figure 17).
+In Feature Concatenation, the average pooling features are combined with the country categorical feature and then passed into the final fully-connected layer.
+Inspired by [21], in Geometric Fusion, features before average pooling are concatenated and processed through a CNN to learn feature importance.
+Next, the resulting weights are then multiplied by the inputs and also concatenated again to produce the output features.
+Figure 16: Feature Concatenation.
+Figure 17: Geometric Fusion.
+For type II, training utilizes all modalities, while inference is based solely on the two top-view inputs.
+Inspired by [22], we adopt a Shared-Specific Feature Modeling approach to address the missing modality issue (Figure 18).
+Each modality is assigned a specific encoder, with street view and VHR inputs that also share an encoder.
+The shared encoder aligns the features of the VHR and street-view.
+During inference, aligned VHR features are used in place of street-view features.
+In type II, two auxiliary losses are added: L1 Loss to align shared features and Cross Entropy Loss to differentiate specific features.
+The street-view and VHR modalities use the same encoder for both shared and specific encoders, while the S2 modality uses a different, shallower encoder.
+We also apply input dropout, randomly replacing street-view images with all-zero images.
+Figure 18: Shared-Specific Feature Modelling for type II, developed by the team xmb, i.e. the second best model.
+VII-D Results and Discussion
+Experimental Design.
+We implement our approaches using the PyTorch library, the PyTorch Lightning framework, and timm [23] library for the feature encoders.
+The efficient and lightweight encoders from timm were chosen for our experiments, with Efficientnetv2_s, Mobilevitv2_150, Efficientnetv2_b3 for type I, and Efficientnetv2_b1 (street-view and VHR), Efficientnetv2_b0 (S2) for type II.
+We train 4 models for type I and use their ensemble as the final model.
+Also, only 2 models are trained for type II.
+For most of the models, we use a batch size of 40.
+The models are trained using the AdamW optimizer (i.e.
+learning rate: 1e-4, weight decay: 1e-3) and a CosineAnnealingLR scheduler (i.e.
+eta min: 5e-5) for 30 epochs without early stopping, applying Focal Loss with Label Smoothing (i.e.
+smoothing coefficient: 0.1).
+The train dataset is split into 10 stratified folds, with each model using one fold for validation and the remaining 9 for training.
+Street view and VHR inputs are resized to 512x512 and normalized to [0,1].
+S2 inputs are clipped at a maximum value of 10,000 and normalized to [0,1].
+Using the Albumentations library, we apply augmentations such as CoarseDropout, GridDropout, and Spatter to enhance the robustness of the street-view input during training.
+We also apply test-time augmentation for both model types using Flip transformations.
+VII-E Detailed Results and Discussion
+Our submission, using the proposed methods, achieves scores of 0.7122 for MPA_ALL_MODALITIES and 0.5814 for MPA_TOP_VIEW on the private leaderboard.
+These results demonstrate the effectiveness of the shared-specific feature modeling approach in addressing the missing modality issue.
+For type I, only the features at the end of the encoders are fused by late fusion, leaving the middle-layer features unconnected.
+Integrating middle fusion could be an interesting avenue for future work, and utilizing attention mechanisms for fusion may also be promising.
+VIII The third best model: The AI Buzzard
+SwinT-encoder late data fusion with attention.
+The model Swin Transformer (SwinT) pretrained on ImageNet22K is used by the team The AI Buzzard, for the encoders of the three different input modalities.
+Late data fusion using attention is performed, as well as 5-fold cross-validation training, grouped by cities and stratified by the class label.
+Data fusion is performed by computing a weighted sum of the embeddings.
+In this way, the model is able to handle missing data during inference (i.e.
+missing street-view images).
+The flowchart diagram is shown in Fig.
+19.
+For this method, the models are first trained on all the data and, then, they are fine-tuned on the individual countries.
+For data augmentation, random flipping, color jittering, and random resized cropping are used.
+Figure 19: Flowchart of the model that uses SwinT encoders, developed by the team The AI Buzzard, i.e. the third best model.
+VIII-A Methodology
+We develop a multi-modal model that can process orthophotos, street-view photos, and Sentinel-2 imagery to predict the building age of houses. The code is available in Github [24].
+VIII-B Data Preparation
+We apply data augmentation.
+All orthophotos are resized to fit the expected resolution of the model.
+We apply random horizontal and vertical flips with a probability of $p=0.5$, as well as color jitter, varying brightness, contrast, and saturation.
+All street-view facade photos are randomly cropped and resized during training, and also resized in inference.
+Sentinel-2 images are processed to fit standard computer vision models.
+Original images ($64\times 64$ pixels, 12 channels) are rearranged to $128\times 128$ pixels and 3 channels: the top left corner contains the NDVI, NDWI, and NDBI; top right B04, B03, B02; bottom left B05, B06, B07; and bottom right B8A, B11, B12 bands, respectively.
+VIII-C Model
+We developed a flexible and robust multi-modal model.
+Each input modality (street-view, orthophoto, or Sentinel-2) is processed by a pretrained SwinV2 encoder [25] from the TIMM collection [26].
+These models were pretrained on ImageNet to extract general image features.
+The resulting embedding vectors are combined by late fusion [27].
+Each embedding is processed through a linear layer that yields a learning attention weight.
+Then, the weighted sum of the embeddings is calculated and passed to a final classification layer.
+Thus, if a modality is missing during inference, the fusion module is robust, and the network still yields a prediction.
+The model can be used as a one-size-fits-all solution to predict building age, and this is important. We note that the results are improved when we train two different models:
+VIII-C1 All three modalities (country ID QCD, HUN)
+We use a SwinV2 transformer in base size for embedding the streetview photos and the orthophotos.
+The transformer has 87.9 million parameters and expects inputs of $384\times 384$ pixels.
+It can take advantage of the high resolution of the streetview and orthophotos.
+VIII-C2 Only top modalities (country ID PNN, FMW)
+We use small SwinV2 transformers (49.7 million parameters, $256\times 256$ pixels) for embedding orthophotos and Sentinel-2.
+VIII-D Results and Discussion
+VIII-D1 Development set
+The dataset contains images from different countries and cities, identified by a country ID and city ID.
+Only four countries are represented in the test dataset (QCD, HUN, FMW, and PNN).
+We split a development set with 1000 samples from the training dataset, following the distribution of country IDs in the test dataset.
+VIII-D2 Cross-Validation
+We perform 5-fold stratified grouped cross validation.
+The dataset is split such that samples with the same city ID are grouped together.
+In addition, the split is stratified such that the class label distribution in each split is close to the overall class label distribution.
+VIII-D3 Training
+We train first with all available data.
+Then, we fine tune the model using samples only from countries QCD, FMW, and PNN, respectively.
+For predicting on HUN samples, we use the model finetuned on QCD samples.
+We use the cross-entropy loss function with class weights to counter class imbalance.
+Early stopping is applied with a patience of six epochs.
+Training is performed on an NVIDIA A100 GPU with 40 GB memory.
+Fine tuning the SwinV2 transformers and training the fusion module takes no more than five epochs.
+VIII-D4 Inference
+We make predictions with the five pretrained models and choose the majority class.
+If there is a tie, then we choose the building age that is more likely based on the distribution in the training data.
+VIII-E Detailed Results and Discussion
+Figure 20: Confusion matrix for samples including all modalities.
+We evaluate the model that is finally submitted to the development set.
+The mean average precision (MAP) is 0.7147 on samples including all modalities, and 0.6013 on samples including only top-view modalities.
+Figure 20 shows the confusion matrix for samples that includes all modalities.
+Class labels are ordered from 0 (oldest buildings) to 6 (newest buildings).
+We note that neighboring classes are often mistaken for each other.
+The network tends to predict the oldest building age (i.e.
+class 0) too often, likely due to the fact that it is the most probable class in the provided multi-modal training dataset.
+Discussion about the key features.
+The main characteristics of all the models discussed in the preceding paragraphs are: i) very good/ top pretrained model, i.e.
+EVA [9] or EVA-02 [10], that is used as a starting point so as to not train from scratch, ii) model ensemble, iii) hyper-parameter optimization using for example Optuna, iv) label correlation modelling between adjacent classes (per modality) because the distance between adjacent classes is smaller than the distance between classes that are not adjacent, and v) missing modality distribution alignment loss together with a shared feature encoder.
+Here, additional key features of the main models are: K-fold cross-validation training, geometric data fusion using the methodology in [19], and use of a categorical feature for the country in the form of one-hot vector representation.
+IX Baseline model: Age of Buildings for Sustainability (ABS)
+We present the baseline model Age of Buildings for Sustainability (ABS) which is based on Transformers, and more specifically on the recently proposed model SegFormer [28, 29], which uses multi-scale features.
+This baseline model of the challenge employs three encoder networks, one for each of the modalities of the dataset MyCD, and performs late data fusion by concatenating the feature representations in the latent space.
+The flowchart diagram of the model ABS is shown in Fig.
+21.
+The output is the construction epoch class for the queried building.
+The model ABS is trained, validated, and evaluated on the multi-modal dataset MyCD.
+For the two encoders for the modalities street-view and top-view VHR, we use the model SegFormer B5.
+Also, for the Sentinel-2 images, we use an EO Foundation Model that has been trained on unlabelled global Sentinel-2 data [30, 31].
+The pre-training task was geo-location prediction and all spectral bands were used.
+The architecture is a geo-aware U-Net-based model.
+The model ABS combines different models trained for each image modality independently.
+We also release our code for ABS for reproducibility [^1] The GitHub repository is: http://github.com/AI4EO/MapYourCity.
+In addition, the multi-modal dataset MapYourCity is in: http://www.eotdl.com/datasets/AI4EO-MapYourCity..
+Figure 21: Flowchart of the model ABS, the baseline of the challenge we organized using the multi-modal dataset MyCD.
+Figure 22: Left: Classification confusion matrix evaluation results for the model ABS for the multi-modal dataset MyCD.
+Private leaderboard of the challenge.
+Right: Confusion matrix for the model ABS described in Sec.
+IX.
+Public leaderboard of challenge.
+Fig.
+22 shows the classification confusion matrix evaluation re- sults for the model ABS for the multi-modal dataset MyCD for the private leaderboard (left) and for the public leaderboard of the challenge (right).
+The evaluation result for the model ABS for cities that are in the training data, i.e.
+In-Distribution (ID), in the MPA metric is $64.61$, while MPA is $60.11$ when the S-2 modality is not used.
+Moreover, when the two top-view modalities (i.e.
+VHR and S-2) are both not utilized, then the MAP evaluation metric is $58.72$.
+X Evaluation and Results
+In this work, for the numerical evaluation results, we compute the classification confusion matrix for the task of predicting the construction period of buildings, where we have seven different classes, for example “After 2006”.
 We calculate the average of the diagonal items of the confusion matrix, and we report this metric, which is also known as the Mean Producer Accuracy (MPA) evaluation metric.
-We present the evaluation of the models from the challenge in Figs.
-11 and 12.
+We present the evaluation results of the models from the challenge in Figs.
+6 and 7.
 We use MPA_ALL_MODALITIES to denote the setting when both training and testing are performed with all modalities.
@@
 In Figs.
-11 and 12, in the results [^2] https://platform.ai4eo.eu/map-your-city/leaderboard, the MPA_TOP_VIEW scores are lower than MPA_ALL_MODALITIES.
-The top score for MPA_ALL_MODALITIES in the private leaderboard is $0.74$ indicating that the estimation of the construction epoch of buildings in unseen cities is possible.
-When using only the top-view data during inference, then the performance of the models decreases by more than $10$ points in the MPA score.
+6 and 7, in the results [^2] https://platform.ai4eo.eu/map-your-city/leaderboard, the MPA_TOP_VIEW scores are lower than MPA_ALL_MODALITIES.
+The top score for MPA_ALL_MODALITIES in the private leaderboard is $0.74$ indicating that the estimation of the construction epoch of buildings in unseen cities is possible using these models.
+Furthermore, when using only the top-view data during inference, then the performance of the models decreases by more than $10$ points in the MPA score.
 Finally, we observe that for the TOTAL scores, the top winning values are approximately $0.65$ in the private leaderboard.
-VI Conclusion
+XI Conclusion
 In this work, we have introduced the new multi-modal benchmark dataset MyCD and the real-world multi-modal application study for EO generalization.
@@
 Acknowledgment
-The challenge was accomplished thanks to the contributions of many people. We want to thank all the team: Juan Pedro, Dennis Albrecht, and everyone in ESA AI4EO.
+The challenge was accomplished thanks to the contributions of many people. We want to thank all the team: Juan Pedro, Dennis Albrecht, and everyone in ESA AI4EO http://ai4eo.eu.